Rusty Nails Bar & Grill reported free cash flow of $757,000.00 last year. Free cash flows are projected to grow by 2.50% annually. For the same period, the firm reported marketable securities of $9,100.00 and $620,000.00 in outstanding long-term debt. The firm’s reported shareholder equity totaled $300,000.00. The firm currently has 100,000.00 common shares outstanding and the cost of capital is 14.00%.

What is the value of operations for Rusty Nails Bar & Grill?

What is the total corporate (firm) value for Rusty Nails Bar & Grill?

What is the intrinsic value of common equity for Rusty Nails Bar & Grill?

What is the intrinsic value of common equity per share for Rusty Nails Bar & Grill?
Answer #1

Value of operations =Free cash Flow*(1+growth)/Cost of Capital-Growth)=757000*(1+2.5%)/(14%-2.5%) =6747173.91
Total Corporate Value = Value of Equity+Marketable securities+Debt =300000+9100+620000 =929100
Intrinsic value = Value of operations -Marketable securities-Debt =6747173.91-9100-620000=6118073.91
Value per share =Intrinsic value/Number of shares =6118073.91/100000 =61.18